OpenStudy (anonymous):

Jim's uncle is 4 times as old as his nephew Jim. In 10 years, Jim's uncle's age will be 20 years more than twice Jim's age. How old is Jim's uncle right now?

OpenStudy (anonymous):

We can write: U = 4J Where 'U' is his uncle's age and 'J' is Jim's age. It says: In 10 years their ages will be: Uncle = J + 10 Jim = U + 10 So we can write: U + 10 = 2(J + 10) + 20 U + 10 = 2J + 40 U - 2J = 30 And we know U = 4J So we plug that in: 4J - 2J = 30 Subtract 4J - 2J @Dustinself42

OpenStudy (anonymous):

Thank you sooo much!!!!!

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Jim's uncle is 4 times as old as his nephew Jim. In 10 years, Jim's uncle age will be 20 years more than twice Jim's age. How old is Jim right now?